the CO2 gas reaches the fire in 3.9 seconds, at what time will the CH4 reach the fire?

Your answer


Submit




1
Expert's answer
2021-06-07T03:23:54-0400

Graham's law of effusion:

"\frac{Rate_1}{Rate_2}= \sqrt{\frac{M_2}{M_1}} \\

Rate_1=\frac{L}{t_1} \\

Rate_2 = \frac{L}{t_2} \\

M_1(CO_2) = 44 \;g/mol \\

M_2(CH_4) = 16 \;g/mol \\

\frac{\frac{L}{t_1}}{\frac{L}{t_2}}= \sqrt{\frac{16}{44}} \\

\frac{t_2}{t_1}= 0.603 \\

t_1=3.9 \;s \\

t_2=0.603t_1 \\

= 0.603 \times 3.9 = 2.35 \;s"

Answer: 2.35 s